(no subject)

Mark Timko (timko@ktev3)
Sun, 20 Jun 1999 07:18:14 -0500 (CDT)

Shift Summary for hhh shift, dd/mmm/99

On Shift: S. Glazov, M. Timko

Hours of beam delivered at all intensities: 7.0
Hours of beam delivered at nominal intensity: 7.0
Hours of E832 data (special runs not included): 7.0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: none

========================================================================
Runs ended this shift: Still on 13668; ended no runs.

Tapes brought to FCC this shift: None

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:
E832 datataking only. No muon run by request of run manager.
Monitoring.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

13668 02:40 csiprobs-lazy kumac gave blown caomparator
id = 2310. P. Shanahan paged and investigated
while we let Julie sleep. He said it's not so
bad; not 100% evident that it has failed; he
can correct for it offline. So, wait for
the day-time people to check it out.

13668 03:25 Onsplit wanted two tapes mounted. So we wrote
to both tapes at once. KTEV1 = QKNL01, and
KTEV2 = QKSM01. S.G. worried for a time, but
it completed just fine.

13668 253 03:45 Main Control room called and wanted to change
the beam steering. The beam positions may
move a little. They promised to put them back.


========================================================================
Follow up Notes on Previous Shift : None

========================================================================
Items to Follow up on Next Shift

sdaq_online needs fixing -- expert was notified.

[end of shift summary]